# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

### v1

@@ 1 Introduction
Figure 1: Identifiability of linear properties. Plots in the left and right dotted squares illustrate the embeddings of two next-token predictors $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta$ that generate the same distribution $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}$ within a set of conditionals distributions $p(y\mid{\bm{\mathrm{x}}})$. Theorem 5 proves a one-to-one correspondence (the dashed orange arrow) between conditional distributions and $\sim_{EL}$-equivalent models (the orange partitions of $\Theta$). This extends a result by Roeder et al. [3] characterizing $\sim_{L}$-equivalent models (green partitions of $\Theta$). Here $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\sim_{EL}}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ while the embedding representations are not equal up to a linear transformation (thus $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}){\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\not\sim_{L}}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$), as shown by how the purple and blue parallelograms in the embeddings of the left model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ get distorted in those of the right model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$. Both models display relational linear probing for the query ${\bm{\mathrm{q}}}$$=$Is the text written in English?”: one can linearly separate the embeddings of textual inputs which, when concatenated with ${\bm{\mathrm{q}}}$, have “yes” as the likeliest next token, from those which have “no”. In Proposition 14, we provide conditions under which all or none of the models in the ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\sim_{EL}}$ equivalence class share the same linear property.
@@
In natural language processing, it is well-established that linear relationships between high-dimensional, real-valued vector representations of textual inputs reflect semantic and syntactic patterns. This was motivated in seminal works [4, 5, 6, 7, 8] and extensively validated in word embedding models [9, 10, 11] as well as modern large language models trained for next-token prediction [12, 13, 2, 14, 15, 16, 17, 18, 19].
@@
This ubiquity is puzzling, as different internal representations can produce identical next-token distributions, resulting in distribution-equivalent but internally distinct models. This raises a key question: Are the observed linear properties shared across all models with the same next-token distribution?
Our main result is a mathematical proof that, under suitable conditions, certain linear properties hold for either all or none of the equivalent models generating a given next-token distribution. We demonstrate this through three main contributions.
@@
The first main contribution (Section 3) is an identifiability result characterizing distribution-equivalent next-token predictors. Our result is a generalization of the main theorems by Roeder et al. [3] and Khemakhem et al. [20], relaxing the assumptions of diversity and equal representation dimensionality. This result is of independent interest for research on identifiable representation learning since our analysis is applicable to several discriminative models beyond next-token prediction [3].
@@
Our second main contribution (Section 4) is to subsume several linear properties in a common framework. We start by defining an analogue to relational linearity [1, 2], where the definition only relies on terms appearing in our identifiability result. The key idea is to represent entities as vectors, binary relations as matrices, and to model the operation of applying a relation to an entity through matrix-vector multiplication, which yields the vector corresponding to the related entity. For example, in the sentence “Jimi Hendrix plays the guitar”, the relation between the entities “Jimi Hendrix” and “the guitar” is signified by the word “plays” and represented as a matrix-vector multiplication in representation space.111“the piano” is a less probable continuation, we only refer to the likeliest one for ease of exposition. We then define relational counterparts to linearity properties described and analyzed in previous works [21, 22, 23, 15, 19], thus making them amenable to our analysis.
@@
Our third main contribution (Section 5) is to show that under suitable conditions, these linear properties either hold in all or none of the models generating a given distribution. For this, we combine the definitions in Section 4 and our characterization of distribution-equivalent next-token prediction models in Section 3. Thus, identifiability theory enables us to explain what linearity properties are shared across language models which are equivalent next-token predictors.
@@ 2 Preliminaries
Next-token predictors. Here, we introduce the general form of next-token predictors used in our analysis. [^2] In Appendix A we show, using a derivation by [3], that decoder-only transformer models reduce to this form. We consider models to take text sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ of tokens $\mathcal{A}$ as input. A next-token predictor $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ consists of two functions: ${\bm{\mathrm{f}}}:\mathsf{Seq}(\mathcal{A})\to\mathbb{R}^{d}$ maps sequences to their representation, called embeddings, and ${\bm{\mathrm{g}}}:\mathcal{A}\to\mathbb{R}^{d}$ maps tokens to their representation, called unembeddings. Let $\Theta_{d}$ be the set of all tuples $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ with representation dimensionality $d$ and let $\Theta:=\bigcup_{d=1}^{\infty}\Theta_{d}$ be the set of all tuples with any dimensionality $d$. A next-token predictor models the conditional distribution of the next-token $x_{t+1}$ given the context ${\bm{\mathrm{x}}}_{1:t}$ as
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(x_{t+1}\mid{\bm{\mathrm{x}}}_{1:t}):=\frac{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1:t})^{\top}{\bm{\mathrm{g}}}(x_{t+1}))}{Z({\bm{\mathrm{x}}}_{1:t})}$ (1)
where $Z({\bm{\mathrm{x}}}_{1:t}):=\sum_{y\in\mathcal{A}}\exp[{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1:t})^{\top}{\bm{\mathrm{g}}}(y)]$ is the normalizing constant. Models of the form of 1 are trained to maximize the conditional log-likelihood of the data. Provided a data distribution $p_{\mathcal{D}}$ over sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$, we consider the log-likelihood:
$\mathcal{L}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}):=\mathbb{E}_{{{\bm{\mathrm{x}}}\sim}p_{\mathcal{D}}}\left[\sum_{t=1}^{T({\bm{\mathrm{x}}})-1}\log p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(x_{t+1}|{\bm{\mathrm{x}}}_{t:1})\right]$ (2)
where $T({\bm{\mathrm{x}}})$ indicates the length of the sequence ${\bm{\mathrm{x}}}$. For a fixed dimension $d$, the learning objective can be written as
@@
Since we work in the non-parameteric setting, the functions ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}$ are unrestricted and, thus, the expressivity of the model is determined only by $d$, the dimensionality of the representation space.
Representation dimensionality and approximation capacity. In theory and for real-valued inputs, models of the form in 1 have been proven to be universal approximators i.e., they can approximate any conditional distribution $p(x_{t+1}\mid{\bm{\mathrm{x}}}_{1:t})$ to arbitrary precision, given a sufficiently large representation dimensionality $d$ [20]; similar results may apply to next-token predictors. In practice, even if a representation dimensionality of $d$ may be sufficient to represent a given distribution well, different practitioners may choose models with representations of different dimensionality. The linear identifiability results by Khemakhem et al. [20], Roeder et al. [3] cannot be applied in this setting, since they consider models with equal representation dimensionality. Our Theorem 5 alleviates this tension between theory and practice: It characterizes identifiability of next-token predictors modeling the same conditional distribution irrespective of their representation dimensionality.
@@ 3 Identifiability of next-token predictors
We introduce a novel identifiability analysis for the model in 1. In general, a statistical model $p_{\bm{\mathrm{\theta}}}({\bm{\mathrm{x}}})$ parameterized by ${\bm{\mathrm{\theta}}}$ is said to be identifiable up to an equivalence relation $\sim$ over parameters if $p_{\bm{\mathrm{\theta}}}=p_{\tilde{\bm{\mathrm{\theta}}}}\implies{\bm{\mathrm{\theta}}}\sim\tilde{\bm{\mathrm{\theta}}}$. In other words, if two parameters entail the same distribution, then they must be equivalent. The precise notion of equivalence depends on the problem setting. Although it is less commonly discussed, this implication can often be shown to hold also in the other direction so that $\sim$ is effectively a characterization of distribution-equivalent models, i.e., $p_{\bm{\mathrm{\theta}}}=p_{\tilde{\bm{\mathrm{\theta}}}}\iff{\bm{\mathrm{\theta}}}\sim\tilde{\bm{\mathrm{\theta}}}$. In this section, we define an equivalence relation over tuples $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ that characterizes models that entail the same next-token prediction distribution. In other words, we want to define an equivalence relation $\sim$ over $\Theta$ such that $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}\iff({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$; we then say $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ are $\sim$-equivalent. Our characterization applies to pairs of models having different dimensionalities, i.e. $d\not=\tilde{d}$, as opposed to previous works [24, 20, 3, 25].
Previous works [24, 20, 3, 25] have shown that, under an assumption on model representations, a linear invertible transformation relates the representations of two distributionally-equivalent models. Intuitively, this assumption, known as variability [24, 20, 25] or diversity condition [3], requires that at least one model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ “uses” the whole space of representations. To state the condition formally, we define the linear space spanned by the image $\text{Im}({\bm{\mathrm{h}}})\subseteq\mathbb{R}^{d}$ of a function ${\bm{\mathrm{h}}}$ as $\mathrm{SIm}({\bm{\mathrm{h}}}):=\mathrm{span}\{{\bm{\mathrm{v}}}\mid{\bm{\mathrm{v}}}\in\mathrm{Im}({\bm{\mathrm{h}}})\}$. Further, for the unembeddings, we fix an arbitrary token $y_{0}\in\mathcal{A}$ to be a pivot token for the rest of the paper and define:
@@
for all tokens $y\in\mathcal{A}$. The diversity condition [20, 3, 25] can then be defined as:
@@
We say a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ with representation dimensionality $d$ satisfies the diversity condition, if $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})=\mathbb{R}^{d}$.
Intuitively, the diversity condition states that the dimension of the spaces spanned by the output of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ match the dimension of the representation space. When both the diversity condition and $d=\tilde{d}$ hold, existing identifiability results for models similar to 1 (presented in Corollary 6) guarantee equivalence of representations up to an invertible linear transformation [20, 3]. Next, we show how to relax these two conditions via a more permissive equivalence relation.

### v2

@@ 1 Introduction
Figure 1: Identifiability of linear properties. Plots in the left and right dotted squares show the embeddings of two next-token predictors $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta$ that generate the same distribution $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}$ within a set of conditionals distributions $p(y\mid{\bm{\mathrm{x}}})$. Theorem 5 proves a one-to-one correspondence (the dashed orange arrow) between conditional distributions and $\sim_{EL}$-equivalent models (the orange partitions of $\Theta$). This extends a result by Roeder et al. [2021] characterizing $\sim_{L}$-equivalent models (green partitions of $\Theta$). Here $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\sim_{EL}}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ while the embedding representations are not equal up to a linear transformation (thus $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}){\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\not\sim_{L}}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$), as shown by how the purple and blue parallelograms in the embeddings of the left model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ get distorted in those of the right model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$. Both models display relational linear probing for the query ${\bm{\mathrm{q}}}$$=$Is the text written in English?”: one can linearly separate the embeddings of textual inputs which, when concatenated with ${\bm{\mathrm{q}}}$, have “yes” as the likeliest next token, from those that yield “no”. In Theorem 14, we provide conditions under which all or none of the models in the ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\sim_{EL}}$ equivalence class share the same linear property.
@@
In natural language processing, it is well-established that linear relationships between high-dimensional, real-valued vector representations of textual inputs reflect semantic and syntactic patterns. This was motivated in seminal works [Rumelhart and Abrahamson, 1973, Hinton et al., 1986a, b, Rumelhart et al., 1986, Bengio et al., 2000] and extensively validated in word embedding models [Mikolov et al., 2013a, b, Pennington et al., 2014] as well as modern large language models trained for next-token prediction [Burns et al., 2022, Merullo et al., 2023, Tigges et al., 2023, Pal et al., 2023, Gurnee and Tegmark, 2023, Bricken et al., 2023].
@@
This ubiquity is puzzling, as different internal representations can produce identical next-token distributions, resulting in distribution-equivalent but internally distinct models. This raises a key question: Are the observed linear properties shared across all models with the same next-token distribution? Our main result is a mathematical proof that, under suitable conditions, certain linear properties hold for either all or none of the equivalent models generating a given next-token distribution. We demonstrate this through three main contributions.
@@
The first main contribution (Section 3) is an identifiability result characterizing distribution-equivalent next-token predictors. Our result is a generalization of the main theorems by Roeder et al. [2021] and Khemakhem et al. [2020a], relaxing the assumptions of diversity and equal representation dimensionality. This result is of independent interest for research on identifiable representation learning since our analysis is applicable to several discriminative models beyond next-token prediction [Roeder et al., 2021].
00footnotetext: $\clubsuit$ Work done while at the University of Copenhagen.00footnotetext: $\spadesuit$ Shared last author.
@@
Our second main contribution (Section 4) is to subsume several linear properties in a common framework. We start by defining an analogue to relational linearity [Paccanaro and Hinton, 2001], where the definition only relies on terms appearing in our identifiability result. The key idea is to represent entities as vectors, binary relations as matrices, and to model the operation of applying a relation to an entity through matrix-vector multiplication, which yields the vector corresponding to the related entity. For example, in the sentence “Jimi Hendrix plays the guitar”, the relation between the entities “Jimi Hendrix” and “the guitar” is signified by the word “plays” and encoded as a matrix-vector multiplication in representation space. We then define relational counterparts to linearity properties described and analyzed in previous works [Arora et al., 2016, Allen and Hospedales, 2019, Park et al., 2024a, Heinzerling and Inui, 2024], thus making them amenable to our analysis.
@@
Our third main contribution (Section 5) is to show that under suitable conditions, these linear properties either hold in all or none of the models generating a given distribution. For this, we combine the definitions in Section 4 and our characterization of distribution-equivalent next-token prediction models in Section 3. Identifiability theory thus enables us to explain what linearity properties are shared across language models which are equivalent next-token predictors. We illustrate this result in Figure 1.
@@ 2 Preliminaries
Next-token predictors. Here, we introduce the general form of next-token predictors used in our analysis. [^1] In Appendix A we show that decoder-only transformer models can be expressed in this form [Roeder et al., 2021]. We consider models which take text sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ of tokens $\mathcal{A}$ as input. A next-token predictor $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ consists of two functions: ${\bm{\mathrm{f}}}:\mathsf{Seq}(\mathcal{A})\to\mathbb{R}^{d}$ maps sequences to their representations, called embeddings, and ${\bm{\mathrm{g}}}:\mathcal{A}\to\mathbb{R}^{d}$ maps tokens to their representations, called unembeddings. Let $\Theta_{d}$ be the set of all tuples $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ with representation dimensionality $d$ and let $\Theta:=\bigcup_{d=1}^{\infty}\Theta_{d}$ be the set of all tuples with any dimensionality $d$. A next-token predictor models the conditional distribution of the next-token $x_{t+1}$ given the context ${\bm{\mathrm{x}}}_{1:t}$ as
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(x_{t+1}\mid{\bm{\mathrm{x}}}_{1:t}):=\frac{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1:t})^{\top}{\bm{\mathrm{g}}}(x_{t+1}))}{Z({\bm{\mathrm{x}}}_{1:t})},$ (1)
where $Z({\bm{\mathrm{x}}}_{1:t}):=\sum_{y\in\mathcal{A}}\exp[{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1:t})^{\top}{\bm{\mathrm{g}}}(y)]$ is a normalizing constant. Models of the form in Equation (1) are trained to maximize the conditional log-likelihood of the data. For a data distribution $p_{\mathcal{D}}$ over sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$, its log-likelihood is given by:
$\mathcal{L}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}):=\mathbb{E}_{{{\bm{\mathrm{x}}}\sim}p_{\mathcal{D}}}\left[\sum_{t=1}^{T({\bm{\mathrm{x}}})-1}\log p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(x_{t+1}|{\bm{\mathrm{x}}}_{1:t})\right],$ (2)
where $T({\bm{\mathrm{x}}})$ indicates the length of the sequence ${\bm{\mathrm{x}}}$. For a fixed representation dimensionality $d$, the next-token prediction objective can be written as
@@
We consider a setting where both ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}$ are nonparametric functions, so the model’s expressivity is determined solely by the parameter $d$, corresponding to the dimensionality of the representation space.
Representation dimensionality and approximation capacity. In theory and for real-valued inputs, models of the form in Equation (1) have been proven to be universal approximators, i.e., they can approximate any conditional distribution $p(x_{t+1}\mid{\bm{\mathrm{x}}}_{1:t})$ to arbitrary precision, given a sufficiently large representation dimensionality $d$ [Khemakhem et al., 2020a]; similar results may apply to next-token predictors. In practice, even if a representation dimensionality of $d$ may be sufficient to represent a given distribution well, different practitioners may choose models with representations of different dimensionality. The linear identifiability results by Khemakhem et al. [2020a], Roeder et al. [2021] cannot be applied in this setting, since they consider models with equal representation dimensionality. Our Theorem 5 alleviates this tension between theory and practice: It characterizes identifiability of next-token predictors modeling the same conditional distribution irrespective of their representation dimensionality.
@@ 3 Identifiability of next-token predictors
We introduce a novel identifiability analysis for the model in Equation (1). In general, a statistical model $p_{\bm{\mathrm{\theta}}}({\bm{\mathrm{x}}})$ parameterized by ${\bm{\mathrm{\theta}}}$ is said to be identifiable up to an equivalence relation $\sim$ in the model class $\Theta$ if $p_{\bm{\mathrm{\theta}}}=p_{\tilde{\bm{\mathrm{\theta}}}}\implies{\bm{\mathrm{\theta}}}\sim\tilde{\bm{\mathrm{\theta}}}$. In other words, if two parametrizations ${\bm{\mathrm{\theta}}},\tilde{\bm{\mathrm{\theta}}}$ yield the same distribution, then they coincide under the equivalence relation $\sim$. The precise notion of equivalence depends on the problem setting. Although it is less commonly discussed, this implication can often be shown to hold also in the other direction so that $\sim$ is effectively a characterization of distribution-equivalent models, i.e., $p_{\bm{\mathrm{\theta}}}=p_{\tilde{\bm{\mathrm{\theta}}}}\iff{\bm{\mathrm{\theta}}}\sim\tilde{\bm{\mathrm{\theta}}}$. In this section, we define an equivalence relation over tuples $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ that characterizes models that entail the same next-token distribution. In other words, we want to define an equivalence relation $\sim$ over $\Theta$ such that $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}\iff({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$; we then say $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ are $\sim$-equivalent. Our characterization applies to pairs of models having different dimensionalities, i.e., $d\not=\tilde{d}$, as opposed to previous works only considering $d=\tilde{d}$ [Khemakhem et al., 2020b, a, Roeder et al., 2021, Lachapelle et al., 2023].
Previous works have shown that, under an assumption known as variability [Khemakhem et al., 2020b, a, Lachapelle et al., 2023] or diversity condition [Roeder et al., 2021], the representations extracted by distributionally-equivalent models are equal up to a linear invertible transformation. Intuitively, requires that at least one model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ “spans” the whole representation space. To formally state the condition, we define the linear space spanned by the image $\text{Im}({\bm{\mathrm{h}}})\subseteq\mathbb{R}^{d}$ of a function ${\bm{\mathrm{h}}}$ as $\mathrm{SIm}({\bm{\mathrm{h}}}):=\mathrm{span}\{{\bm{\mathrm{v}}}\mid{\bm{\mathrm{v}}}\in\mathrm{Im}({\bm{\mathrm{h}}})\}$. Additionally, for the unembeddings, we choose an arbitrary token $y_{0}\in\mathcal{A}$ as a pivot for the remainder of the paper and define:
@@
for all tokens $y\in\mathcal{A}$. The diversity condition can then be defined as follows:
@@
We say that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ with representation dimensionality $d$ satisfies the diversity condition if $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})=\mathbb{R}^{d}$.
Intuitively, the diversity condition states that the dimension of the spaces spanned by the output of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ match the dimension of the representation space. When both the diversity condition and $d=\tilde{d}$ hold, existing identifiability results for models of the form in Equation (1) (presented in Corollary 6) guarantee equivalence of representations up to an invertible linear transformation [Khemakhem et al., 2020a, Roeder et al., 2021]. Next, we show how to relax these two conditions via a more permissive equivalence relation.
Figure 2: Illustration of the $\sim_{EL}$ equivalence relation. (Left) In the leftmost model, $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, the embeddings lie on a manifold $\mathrm{Im}({\bm{\mathrm{f}}})\subsetneq\mathbb{R}^{3}$, yielding $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathbb{R}^{3}$. To ease visualization, $\mathrm{Im}({\bm{\mathrm{f}}})$ is plotted as a continuous manifold in the figure, although in practice textual inputs are discrete. The unembeddings lie on a two-dimensional space, $\mathrm{SIm}({\bm{\mathrm{g}}}_{0})\cong\mathbb{R}^{2}$, drawn in light blue. Consequently the projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ map onto a two-dimensional subspace, i.e., ${\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{N}}={\bm{\mathrm{P}}}_{\mathcal{G}}$. (Right) The rightmost model, $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, represents both the embeddings and the unembeddings in a two-dimensional space. We therefore have $\mathrm{SIm}(\tilde{{\bm{\mathrm{f}}}})=\mathrm{SIm}(\tilde{{\bm{\mathrm{g}}}_{0}})=\mathbb{R}^{2}$, which implies that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}={\bm{\mathrm{I}}}$. Thus applying these projection matrices to embeddings and unembeddings leaves them unchaged (top-right and bottom-right grids). (Center) The equivalence relation $\sim_{EL}$ specifies that both ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}$, as well as ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}\tilde{{\bm{\mathrm{g}}}}$, are related by linear invertible transformations defined by the matrices ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}\in\mathbb{R}^{3\times 2}$.